OpenStudy (imtiaz7):

what is the potential energy of of a charge of 2.5*10^-6 C at a point 150 mm from the centre of sphere?

OpenStudy (festinger):

What is the charge distribution like? Is the charge at the center? What is there sphere there for?

OpenStudy (imtiaz7):

there is a single chare which is in the sphere.

OpenStudy (festinger):

use gauss's law to find E, then use this E to find V \[\iint\limits_{Closed\:Surface} E \:dA=\frac{Q_{enclosed}}{\epsilon_{0}}\] \[E=-\frac{dV}{dr}\] Since we're dealing with a sphere, it's symmetrical. So E is the same throughout and is a constant \[E\iint\limits_{Closed\:Surface} \:dA=E4\pi r^{3} = \frac{Q_{enclosed}}{\epsilon_{0}}\] And thus \[E=\frac{Q}{4\pi \epsilon_{0} r^2}\] Since \[E=-\frac{dV}{dr}\] \[Edr=-dV\] and thus \[\int_{\infty}^{r} Edr=\int_{0}^{V} -dV\] \[\int_{\infty}^{r}\frac{Q}{4\pi \epsilon_{0} r^2}dr=-(V-0)\] \[-\frac{Q}{4\pi \epsilon_{0} r}+0=-V\] \[V= \frac{Q}{4\pi \epsilon_{0} r}\] Plug in the relevant values of Q and r and you are done!

OpenStudy (imtiaz7):

by putting values in given eq. i got wrong answer

OpenStudy (imtiaz7):

the answer in my book is 6.375*10^-4 J

OpenStudy (festinger):

So what is the charge of the other particle? Since it's potential energy, it must be between 2 charges